### v1

# Gravitational waves in minimal Maxwell $f(R)$ gravity

Abstract
We discuss the conversion mechanism from scalar field to gravitational waves in magnetic background in $f(R)$ gravity minimal coupled to the Maxwell electrodynamics. Applying the conversion in early universe with primordial magnetic field and the neutron star with strong magnetosphere, the generated gravitational waves, converted from the scalar field, exhibit distinct power spectra which can be potentially probed by future experiments such as DECIGO, BBO, CE and ET.
I Introduction
In last decades, a wide range of cosmological observations have revealed the accelerated expansion of the universe, both in its primordial and late stages. These phenomena challenges General Relativity (GR) in explaining the cosmos evolution without invoking unknown components such as new scalar particles capable of driving inflation at UV scales and dark energy at IR scales. The fact that, till now, no new fundamental particle component has been clearly revealed motivated the exploration of various extended or alternative theories of gravity that, while reproducing the good results of GR, should cure shortcomings at UV and IR scales. See [1] for a comprehensive discussion. In particular, $f(R)$ gravity is a straightforward extension generalizing the Einstein-Hilbert action, linear in the Ricci scalar $R$, to a generic function of it. This approach preserves the geometric foundations of GR while introducing new degrees of freedom and dynamics capable of addressing many cosmic problems to some extents (see for example [2, 3, 4, 5, 6, 7, 8]).
Specifically, it can be demonstrated that $f(R)$ gravity can be recast as a scalar-tensor theory, equipped with a massive scalar and two massless tensor degrees of freedom [9, 10, 11, 12, 13]. The scalar mode corresponds to the scalar perturbation related to the form of $f(R)$ function while the tensor modes are the traceless components which constitute, essentially, the gravitational waves (GWs) of GR. In vacuum, scalar and tensor fields evolve independently in the linear order perturbations, whereas they are coupled to each other in higher order regimes. In particular, the first order of scalar perturbation can induce the second order GWs [14, 15, 16]. Another possibility to couple the scalar and tensor fields is by the inclusion of the matter. One of the simplest choice to introduce matter into dynamics is considering the Maxwell electrodynamics, which is also inspired by the multi-messenger astronomy involving GWs and electromagnetic waves. Incorporating Maxwell electrodynamics into $f(R)$ gravity is explored, for example, in Refs. [17, 18, 19, 20].
In this paper, we find that including the Maxwell electrodynamics, even minimally coupled to $f(R)$ gravity, can source GWs and give rise to a clear signature in the power spectrum. As a starting case, we can consider a pure constant magnetic field background with strength $B$. In addition, we can expand the metric perturbation in a Minkowski background. At linear order in perturbations, the scalar field $\varphi$ can excite GWs via the term $G\varphi B^{2}$, where $G$ is the Newton constant. In the language of particle physics, such a process can be interpreted as a scalaron (scalar perturbation) propagating through the magnetic field that converts into a graviton (tensor perturbation). The process is similar to the graviton-photon and axion-photon conversion in presence of a magnetic background as reported in Refs. [21, 22, 23, 24].
Here we propose the scalaron-graviton conversion in a magnetic field in the framework of a minimal Maxwell $f(R)$ model. However, the probability of conversion from scalaron to graviton is much suppressed due to the presence of $G$ in the source term $G\varphi B^{2}$. Thus we need to search for strong magnetic regimes to enhance the conversion process. There are two interesting potential situations that can be taken into account: the early universe fulfilled with a primordial magnetic field quadratically scaling in redshift [25]; neutron stars with strong magnetosphere where the current strongest record reach a $10^{15}$s level [26]. For the same reason of $G$ suppression, the graviton-photon conversion, in strong magnetic environments, has been explored in Refs. [27, 28, 29, 30, 31].
To estimate the generated GW spectrum, understanding the properties of scalaron is essential. The scalaron mass is constrained in both early and late universe: for example, the Starobinsky inflation fixes $m_{a}\simeq 3\times 10^{13}$eV [32]; the BBN requires $10^{-16}\textrm{eV}<m_{a}<10^{4}\textrm{eV}$ [33], the observed GW speed from binary systems gives $m_{a}<10^{-22}$eV [34] and so on. Some of these mass ranges seem contradictory to each other and hence demands some mechanism to circumvent it, e.g. the chameleon dressing or a time-varying mass, which can be realized in various $f(R)$ models [35]. Since the main purpose of this paper focuses on the scalaron-graviton conversion, we perform a phenomenological approach taking $m_{a}$ to be a free parameter in early universe analysis and sufficient small in neutron star analysis. Regarding the density spectrum of the scalaron, we simply regard it as a radiation (at least during the conversion process) constrained by extra effective relativistic species. Assuming a scale invariant spectrum of scalaron as the simplest scenario, it is possible to show that the generated GWs can be detected by DECIGO [36], BBO [37], CE [38] and ET [39].
The paper is organized as follows. the first order perturbation of Maxwell $f(R)$ gravity is given in Sec. II. The scalaron-graviton conversion probability in constant magnetic background is derived in Sec. III. Two applications in early universe and neutron star are studied in Sec. IV. Discussion and conclusion are reported in Sec. V.
II Maxwell $f(R)$ gravity
Let us consider a generic $f(R)$ model minimally coupled to the Maxwell electrodynamics with the action given by
$S=\int\sqrt{-g}d^{4}x\left(\frac{1}{\kappa^{2}}f(R)-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}\right),$ (1)
where $\kappa=\sqrt{16\pi G}$ and the electrodynamics tensor is $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ with the potential vector $A_{\mu}$. The field equations for metric $g_{\mu\nu}$ can be obtained from the variation of the action:
$R_{\mu\nu}f_{R}-\frac{1}{2}g_{\mu\nu}f(R)+g_{\mu\nu}\nabla^{\alpha}\nabla_{\alpha}f_{R}-\nabla_{\mu}\nabla_{\nu}f_{R}$ $=$ $\frac{\kappa^{2}}{2}\left(g^{\alpha\beta}F_{\alpha\mu}F_{\beta\nu}-\frac{1}{4}g_{\mu\nu}F_{\alpha\beta}F^{\alpha\beta}\right),$ (2)
where $f_{R}=df(R)/dR$ and $\nabla_{\mu}$ is the covariant derivative. The extra scalar mode in $f(R)$ theory can be extracted from the trace of equation
$3\nabla^{\mu}\nabla_{\mu}f_{R}+Rf_{R}-2f(R)=0\,$ (3)
that, with respect to GR, is dynamical. Note that the RHS of above equation vanishes due to the conformal property of the Maxwell electrodynamics. With the definition $\Phi=f_{R}$ and $\frac{dV}{d\Phi}=\frac{2f(R)-Rf_{R}}{3}$, Eq. 3 can be rewritten in form of the Klein Gordon-type equation
$\nabla^{\mu}\nabla_{\mu}\Phi-\frac{dV}{d\Phi}=0.$ (4)
To linearize the theory, we perturb the scalar field as $\Phi=\Phi_{0}+\varphi$ around its ground value $\Phi_{0}$ determined by the minimum of $\left.\frac{dV}{d\Phi}\right|_{\varPhi=\Phi_{0}}=0$. Moreover, we assume the background spacetime is Minkowski and perturb the metric field $g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}$. Thus the equation of $\varphi$ becomes
$\square\varphi-m_{S}^{2}\varphi=0,$ (5)
where $\square=\eta^{\eta\nu}\partial_{\mu}\partial_{\nu}$ is d’Alembert operator and its mass is given by $m_{S}^{2}=\left.\frac{d^{2}V}{d\Phi^{2}}\right|_{\varPhi=\Phi_{0}}$.
Regarding the perturbation of field equations in $f(R)$ theory, it is worth introducing the variable
$\overline{h}_{\mu\nu}=h_{\mu\nu}-\frac{1}{2}\eta_{\mu\nu}h-\eta_{\mu\nu}\varphi,$ (6)
where $h=\eta^{\mu\nu}h_{\mu\nu}$. As demonstrated in Ref. [9, 11], $\overline{h}_{\mu\nu}$ describes two massless tensor/polarizations of GWs as in GR, while $\varphi$ is a massive scalar field arising from the higher-order of metric $f(R)$ gravity [6]. In particle language, perturbations of $\overline{h}_{\mu\nu}$ and $\varphi$ fields are dubbed “graviton” and “scalaron”, respectively. Due to the traceless stress-energy tensor of the Maxwell electrodynamics, we are allowed to impose the Lorenz traceless gauge fixing on $\overline{h}_{\mu\nu}$:
$\partial^{\mu}\overline{h}_{\mu\nu}=0,\qquad\eta^{\mu\nu}\overline{h}_{\mu\nu}=0,$ (7)
and expand the Ricci tensor and LHS of Eq. 2 to the linear order
$R_{\mu\nu}$ $=$ $\frac{1}{2}\left(-\square\overline{h}{}_{\mu\nu}+2\partial_{\mu}\partial_{\nu}\varphi+\eta_{\mu\nu}\square\varphi\right),$
$LHS$ $=$ $R_{\mu\nu}\Phi_{0}-\frac{1}{2}g_{\mu\nu}\left(f(0)+f_{R}(0)R+o(R^{2})\right)+\eta_{\mu\nu}\square\varphi-\partial_{\mu}\partial_{\nu}\varphi$ (8)
$=$ $-\Phi_{0}\frac{1}{2}\eta^{\alpha\beta}\partial_{\alpha}\partial_{\beta}\overline{h}_{\mu\nu}+\left(\Phi_{0}-1\right)\left(\partial_{\mu}\partial_{\nu}\varphi-\eta_{\mu\nu}\square\varphi\right)$
$=$ $-\frac{1}{2}\square\overline{h}_{\mu\nu}.$
In the above derivation, the $f(R)$ model is required to satisfy the conditions $\Phi_{0}=f_{R}(0)=1$ and $f(0)=0$. A typical example is the Starobinsky model $f=R+\alpha R^{2}$. Finally we obtain the linearized field equations for massless tensor and massive scalar modes of GWs
$\square\overline{h}_{\mu\nu}$ $=$ $\kappa^{2}\left(\overline{h}^{\alpha\beta}F_{\alpha\mu}F_{\beta\nu}-\frac{1}{2}\eta_{\mu\nu}\overline{h}^{\rho\sigma}F_{\alpha\rho}F_{\;\sigma}^{\alpha}+\frac{1}{4}\overline{h}_{\mu\nu}F_{\alpha\beta}F^{\alpha\beta}-\varphi\left(F_{\alpha\mu}F_{\;\nu}^{\alpha}-\frac{1}{4}\eta_{\mu\nu}F_{\alpha\beta}F^{\alpha\beta}\right)\right),$ (9)
$\square\varphi$ $=$ $m_{S}^{2}\varphi.$ (10)
Here the index are raised or lowered by $\eta_{\mu\nu}$ and, from now on, we follow this rule unless explicitly stated. It is worth mentioning that the electrodynamics equation is ignored here because we simply assume a static constant magnetic field. We will expand to include the dynamics of electromagnetic field (i.e. photons) in a forthcoming paper.
III Constant magnetic background
From Eq. 9, we find that the scalar field interacting with the electromagnetic one can source the GWs. In fact, the main purpose of this paper is to show that the scalaron propagating through a magnetic region can convert into gravitons. In order to demonstrate this statement, let us consider scalaron and graviton propagating in a Minkowski spacetime along the $Z$-direction, denoted as $\varphi(t,Z)$ and $\overline{h}_{\mu\nu}(t,Z)$, in presence of a static constant magnetic background. The direction of the magnetic field is assumed to be fixed in the transverse $X$-$Y$plane, for instance along $Y$-direction, namely $\bm{B}=(0,B,0)$. In the transverse traceless gauge with $\overline{h}_{0\mu}=0$, $\partial_{i}\overline{h}^{i}_{\,j}=0$ and $\overline{h}^{i}_{\,i}=0$  [^1] Rigorously, transverse traceless (TT) gauge of $\overline{h}_{\mu\nu}$ can always be imposed in vacuum but not in general cases with non-vanishing sources. The traceless stress-energy tensor of the Maxwell electrodynamics allows a gauge fixing at most to Eq. 7. Using the TT gauge can be regarded as a projection onto the tensorial components of $\overline{h}_{\mu\nu}$ whereas the others are ignored. See Refs. [22, 40, 41, 27] for comments on the TT gauge in relevant topics. , the graviton can be decomposed into plus and cross polarizations as $\overline{h}_{ij}=\overline{h}_{+}e_{ij}^{+}+\overline{h}_{\times}e_{ij}^{\times}$, where bases are $e_{ij}^{+}=X_{i}X_{j}-Y_{i}Y_{j}$ and $e_{ij}^{\times}=X_{i}Y_{j}+X_{j}Y_{i}$. Projecting $e_{ij}^{+}$ and $e_{ij}^{\times}$ on Eqs. 9-10 and using $F_{0\nu}=0,F_{ij}=\varepsilon^{ijk}B_{k}$, we obtain
$\left(-\partial_{t}^{2}+\partial_{Z}^{2}\right)\overline{h}_{+}(t,Z)-\frac{1}{2}\kappa^{2}B^{2}\overline{h}_{+}(t,Z)$ $=$ $-\frac{1}{2}\kappa^{2}B^{2}\varphi(t,Z),$ (11)
$\left(-\partial_{t}^{2}+\partial_{Z}^{2}\right)\overline{h}_{\times}(t,Z)-\frac{1}{2}\kappa^{2}B^{2}\overline{h}_{\times}(t,Z)$ $=$ $0,$ (12)
$\left(-\partial_{t}^{2}+\partial_{Z}^{2}\right)\varphi(t,Z)-m_{S}^{2}\varphi(t,Z)$ $=$ $0$ (13)
The presence/absence of the source term on the RHS of the first two equations is caused by the particular choice of the magnetic field direction, leading to different dynamics of the two polarizations. Since the magnetic effect on polarization is out of the scope of this paper, thus we simply ignore the $\overline{h}_{\times}$ term and denote $\overline{h}_{+}$ as $\overline{h}$ (not the trace of $\overline{h}_{\mu\nu}$).
The term $\kappa^{2}B^{2}\overline{h}$ gives rise to an effective mass of the graviton, $m_{GW}=\kappa B/\sqrt{2}$. We work in the relativistic regime where the frequency of particles is much larger than their mass, namely $\omega\gg m_{GW}$ and $\omega\gg m_{S}$. Performing the Fourier transformation
[…]
and applying the Slowly Varying Envelop Approximation [22] (valid for $\omega\gg\kappa B/2$), $\left(-\partial_{t}^{2}+\partial_{Z}^{2}\right)=\left(\omega^{2}+\partial_{Z}^{2}\right)=2\omega\left(\omega+i\partial_{Z}\right)$, we reduce Eqs. 11 and 13 to the linear differential equations
$i\partial_{Z}\widetilde{h}(\omega,Z)+\left(\omega-\frac{m_{GW}^{2}}{2\omega}\right)\widetilde{h}(\omega,Z)$ $=$ $-\frac{\kappa^{2}B^{2}}{4\omega}\widetilde{\varphi}(\omega,Z),$ (15)
$i\partial_{Z}\widetilde{\varphi}(\omega,Z)+\left(\omega-\frac{m_{S}^{2}}{2\omega}\right)\widetilde{\varphi}(\omega,Z)$ $=$ $0$ (16)
For the boundary condition $\widetilde{h}(\omega,Z_{i})=0$, the solution is
$\widetilde{h}(\omega,Z)=e^{i\omega d}\left(1-e^{-i\frac{\kappa^{2}B^{2}}{4\omega}d}\right)\widetilde{\varphi}(\omega,Z_{i}),$ (17)
where $d=Z-Z_{i}$ is the traveling distance. The conversion rate for scalarons converting into gravitons at a distance $d$ is
$\mathcal{P}_{S\rightarrow GW}(d)=\frac{\left|\widetilde{h}(\omega,Z)\right|^{2}}{\left|\widetilde{\varphi}(\omega,Z_{i})\right|^{2}}=4\textrm{sin}^{2}\left(\frac{\kappa^{2}B^{2}}{8\omega}d\right).$ (18)
Assuming the initial energy density of the scalar source as $\Omega_{\varphi}(\omega,Z_{i})$, we can estimate the abundance of GWs converted from the scalarons in presence of a magnetic field as
$\Omega_{GW}(\omega,Z)=\Omega_{\varphi}(\omega,Z_{i})\mathcal{\mathcal{P}}_{S\rightarrow GW}.$ (19)
We stress that $\mathcal{P}$ is properly called “conversion rate” rather than “conversion probability”, because such a conversion process is unidirectional only from scalarons to gravitons. Thus it can be larger than unity, reaching to maximum 4 from the above expression. Nevertheless, in most physical circumstances, $\mathcal{P}$ is much smaller than the order of unity. Such a suppressed conversion rate leads to a negligible effect on the scalar source, and hence allows one to use always the Klein Gordon equation to describe the scalaron as a free field. Therefore, we simply interpret $\mathcal{P}$ as the conversion probability converting from scalarons to gravitons when $\mathcal{P}\ll 1$, and express Eq. 18 as
$\mathcal{P}_{S\rightarrow GW}=\frac{\kappa^{4}B^{4}d^{2}}{16\omega^{2}}.$ (20)
The conversion mechanism from scalarons to gravitons proposed in this paper is reminiscent of other mixing systems widely studied in literature, for instance, the photon-graviton mixing in magnetic field. Ignoring the plasma effect on photons, the probability of photon-graviton conversion is $\mathcal{P}=\textrm{sin}^{2}(\kappa Bd)$, which reduces to $\mathcal{P}=\kappa^{2}B^{2}d^{2}$ when $\mathcal{P}\ll 1$ [23]. Comparing with Eq. 20, the scalaron-graviton conversion has a quartic dependence on the magnetic field rather than the conventional quadratic standard of the photon-graviton case. Moreover, the extra factor $\kappa^{2}B^{2}/16\omega^{2}<1$ indicates a suppressed probability of scalaron-graviton conversion than the photon-graviton case.
IV Applications
[…]
In most literature, various $f(R)$ models are constructed to explain the inflation or the dark energy, meanwhile the model parameters are constrained by the observational data. As mentioned in Introduction, our main purpose is to take into account the conversion mechanism from scalarons to gravitons in a magnetic background, thus we adopt a phenomenological approach to study its potential effects. We assume the existence of scalarons arisen from $f(R)$ gravity models compatible with Eq. 8 and magnetic field in the whole post-inflation era. Moreover, the scalaron mass $m_{S}$ is taken to be a free parameter. The magnetic field is expected to be scaled in high redshift $z$ as $B=B_{0}z^{2}$, where $B_{0}=\textrm{nGs}$ is the strength of the intergalactic magnetic field today [25]. Especially, it is well within the bound from the BBN constrains $B_{\textrm{BBN}}<2\times 10^{11}\textrm{Gs}$ [42]. Within one Hubble time, the traveling distance is $d=H^{-1}$ and hence conversion probability is
$\mathcal{P}_{S\rightarrow GW}(z)=\frac{1}{16}\left(\frac{\kappa B}{\omega}\right)^{2}\left(\frac{\kappa B_{0}}{H_{0}}\right)^{2},\,\,\,\,\,\,\,\,\omega\gg\kappa B/\sqrt{2},$ (21)
where the frequency scales as $\omega=\omega_{0}z$ and Hubble parameter $H=H_{0}z^{2}$ in R.D. era. In addition, the above formula is only valid at $\omega\gg m_{S}$. With these conditions, the conversion probability reaches a maximum $\mathcal{P}\simeq 6\times 10^{-11}$ at $\omega\simeq\kappa B$. To estimate the generated GW energy density, we consider the simplest case with a scale invariant spectrum of the scalar source. The scalaron can contribute as a radiation component, constrained by extra effective relativistic species $\Delta N_{\textrm{eff}}$ with $\Omega_{\varphi}/(0.2\Omega_{\gamma})<\Delta N_{\textrm{eff}}$  [^2] For a sufficient large mass, scalaron becomes cold DM below a certain temperature in the early universe. Its density abundance can be larger than the radiation one, but there may exist other constrains from the cold DM sector. The different mass cases in left panel of Fig. 1 serves more as a reference.. The Planck constraint $\Delta N_{\textrm{eff}}<0.3$ gives rise to the bound $\Omega_{\varphi}(f)\simeq\Omega_{\varphi}<0.3\times 10^{-5}$ [43]. Taking $\Omega_{\varphi}(f)$ at the constraint bound, we show several cases of different scalaron masses in the left panel of Fig. 1. As a result, the generated GWs can be present in scale invariant power spectrum with energy density $\Omega_{GW}(f)\simeq 10^{-16}$, which could be probed by the DECIGO and BBO. Note that the condition $\omega\gg m_{S}$ constrains the frequency region of the spectrum to $f\gtrsim\sqrt{\kappa B_{0}m_{S}}/2\pi$. For lower frequency regions, Eq. 18 fails and one needs to solve full equation set Eqs. 11 and 13. Nevertheless, we anticipate a damping effect for $\omega<m_{S}$ because it might correspond to the evanescent mode of a localized field in potentials or interactions.
Figure 1: The power spectrum of GWs converted from the scalaron propagating in the primordial magnetic field in early universe (Left) and the strong magnetic field generated by magnetars (Right). The scalaron source is assumed to be in a scale invariant spectrum with density $\Omega_{\varphi}=0.3\times 10^{-5}$. The converting GW spectra are compared to various experiments with their sensitivity curves obtained from [44]. The blue shade in right panel denotes the maximum of GW density by varying magnetic strength of the magnetar and effective magnetosphere size. The damping effect is expected to be dominant when $f\lesssim\sqrt{\kappa B_{0}m_{S}}/2\pi$ (Left) or $f\lesssim\kappa B/2\pi$ (Right), which is schematically shown in dashed lines.
The other potential possibility to enhance the conversion process happens in the magnetosphere of the magnetar, a type of neutron star with very strong magnetic field, where the magnetic field can reach up to $B=10^{13-15}$s [29]. In addition, the size of magnetosphere, $\sim 10^{5}(P/\textrm{s})\textrm{km}$ with spin periods $P$, is generally much larger than the star radius ($\sim 10$km) and the magnetic strength decays with distance as $\sim r^{-3}$ [28, 30]. To avoid the complexity and perform an order-of-magnitude estimation, we effectively model scalarons passing through the magnetosphere as a path with constant magnetic field $B$ and length $d$. From Eq. 20, the conversion probability, presenting in a power law scaling $\omega^{-2}$, indicates a maximum at $\omega\simeq\kappa B$ in the region $\omega\gtrsim\kappa B$. Again, we consider a scale invariant spectrum of the scalaron $\Omega_{\varphi}(f)=0.3\times 10^{-5}$ but its mass is smaller than the frequency of interests $m_{S}<10^{-17}$eV. In a reasonable range of $B<10^{16}\textrm{Gs}$ and $10\textrm{km}<d<100\textrm{km}$, the maximum energy density of GWs converted from scalarons locates inside the blue shade region in the right panel of Fig. 1. The profiles of the GW spectrum are shown in three benchmarks with $B=10^{13,14,15}$s and $d=20$km. The generated GWs can possess energy density in the range $\Omega_{GW}\sim 10^{-16}-10^{-12}$ around frequency region $f\sim 0.1-10$Hz, which can be probed by the future experiments, for instance DECIGO, BBO, ET and CE. We expect a possible damping feature in the low frequency region $\omega<\kappa B$ (dashed lines) caused by the GW scattering in the strong magnetic medium, while this statement needs further detailed analysis.
Finally, two comments are in order. Firstly, as addressed in Ref. [45], it is justified to consider GW propagating through magnetic medium in the flat spacetime background under two conditions $\kappa B<\omega$ and $\kappa BL<1$, where $L$ is the size of the magnetic region. The latter means that the background curvature, sourced by the magnetic field, can be ignored. These conditions are generally satisfied in our model applications in early universe and magnetar. Secondly, the above analysis is based on a scale invariant spectrum of scalaron source and can be easily extended to other forms of spectrum. For instance, one can consider a spectrum peaked at $\omega_{p}$ with a narrow frequency band $\Delta\omega$: the scalaron density is allowed to increase to $\sim\Omega_{\gamma}\omega/\Delta\omega$ in the peaked band. Accordingly, the abundance of generated GWs can be enhanced by a factor $\sim\omega/\Delta\omega$ in the peaked band, rendering its detectable well within the experimental sensitivities.
V Conclusion
In this paper we considered $f(R)$ gravity minimal coupled to the Maxwell electrodynamics, where the system contains a massive scalar field (scalaron), a massless tensor field (graviton/GW) and an electromagnetic field (photon). As a first approximation, we assumed the electromagnetic field to be a constant magnetic field as a background, and focused on the dynamics of the scalaron and the graviton in the Minkowski spacetime. We demonstrated that scalarons propagating in a magnetic field can convert into gravitons. This conversion mechanism is analogous to the graviton-photon, axion-photon conversion widely studied in literature. However, there are two distinct features: the conversion is unidirectional, thus proper to call it “scalaron-induced GW”; the conversion probability has a quartic dependence on the magnetic field. We perform a phenomenological approach to apply this mechanism in the early universe with a primordial magnetic field and the neutron star with a strong magnetosphere. In both cases, the generated GWs in a broad range of parameters can be probed by future experiments (see Fig. 1).
Several interesting aspects can be explored in future researches. An immediate generalization is to include the dynamical electromagnetic fields as photons and plasma effects. Besides, one can employ the techniques handling source-induced GWs in literature to solve full second-order differential Eqs. 11-13 and find out the spectrum for $\omega<\kappa B$. Moreover, dark energy effects could be considered adopting suitable $f(R)$ models and electromagnetic fields. Finally, a de-Sitter background and the chameleon effect should be assumed in this research line [46, 47, 48].
Acknowledgments. We thank Rome Samanta, Andrea Addazi, Gaetano Lambiase and Luca Visinelli for useful discussions. We acknowledge the support of INFN sez. di Napoli, iniziative specifiche QGSKY and MOONLIGHT2. S.C. thanks the Gruppo Nazionale di Fisica Matematica of Istituto Nazionale di Alta Matematica for the support. This paper is based upon work from COST Action CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), supported by COST (European Cooperation in Science and Technology).